### v1

# Revisiting Few-Shot Object Detection with Vision-Language Models

Abstract
Few-shot object detection (FSOD) benchmarks have advanced techniques for detecting new categories with limited annotations. Existing benchmarks repurpose well-established datasets like COCO by partitioning categories into base and novel classes for pre-training and fine-tuning respectively. However, these benchmarks do not reflect how FSOD is deployed in practice. Rather than only pre-training on a small number of base categories, we argue that it is more practical to fine-tune a foundation model (e.g., a vision-language model (VLM) pre-trained on web-scale data) for a target domain. Surprisingly, we find that zero-shot inference from VLMs like GroundingDINO significantly outperforms the state-of-the-art (48.3 vs. 33.1 AP) on COCO. However, such zero-shot models can still be misaligned to target concepts of interest. For example, trailers on the web may be different from trailers in the context of autonomous vehicles. In this work, we propose Foundational FSOD, a new benchmark protocol that evaluates detectors pre-trained on any external datasets and fine-tuned on $K$-shots per target class. Further, we note that current FSOD benchmarks are actually federated datasets containing exhaustive annotations for each category on a subset of the data. We leverage this insight to propose simple strategies for fine-tuning VLMs with federated losses. We demonstrate the effectiveness of our approach on LVIS and nuImages, improving over prior work by 5.9 AP.
Figure 1: We propose a new few-shot object detection (FSOD) benchmark which embraces the challenge of aligning multimodal foundation models to target concepts with vision and language. On the left, we describe the standard setup: methods pre-train on base classes (with many examples per class) and then fine-tune on $K$-shots of novel (and optionally base) classes. On the right, we describe our proposed setup: Given the scale and often private nature of data used to train VLMs, it is impractical to maintain a split of base and novel classes. Instead, one should directly fine-tune VLMs on $K$-shots of the target classes (and evaluate only those target classes). Importantly, VLMs allow us to exploit additional language cues such as class names and descriptions for fine-tuning. We show that such “zero-shot” language cues already outperforms state-of-the-art FSOD methods without any fine-tuning.
Figure 2: Poor Alignment Between VLM and Target Concepts. Although VLMs show impressive zero-shot performance, they struggle when the target class is different from concepts encountered in web-scale training. On the left, we see that the nuImages dataset defines the cab of the truck as a separate concept from its trailer (shown in red). In contrast, the VLM predicts the entire vehicle as a truck (shown in green). On the right, we present the actual class definitions given to the nuImages annotators, provided as both textual descriptions and visual examples. Just as human annotators learn concepts from few-shot multi-modal examples, we argue that VLMs should be fine-tuned with $K$ vision-language examples.
1 Introduction
Object detection is a fundamental problem in computer vision [8, 21] that has matured in recent years [32, 24, 31, 22]. Given a large-scale annotated dataset, one can train a detector from scratch. However, training object detectors for domains with limited annotated data remains challenging, motivating the problem of few-shot object detection.
Status Quo. Few-shot object detection (FSOD) benchmarks have made considerable progress on learning to detect new categories from limited training data. Existing benchmarks are constructed by partitioning popular object detection datasets like PASCAL VOC [5] and COCO [21] into base categories (with many examples per class) and target novel categories (with few examples per class). Detectors are first pre-trained on base classes and are fine-tuned on $K$ examples (or $K$-shots) from novel classes.
Conventional FSOD benchmarks enforce base and novel classes to be disjoint to prevent concept leakage and measure generalization to unseen categories. However, as most detectors are pre-trained on ImageNet, concept leakage already occurs in contemporary benchmarks. For example, cat and person are considered novel in the COCO FSOD benchmark but are already present in ImageNet. Similarly, car is considered novel even though similar concepts like sports car and race car are present in ImageNet. Since concept leakage is difficult to avoid, we argue that it should instead be embraced. Intuitively, pre-training on large-scale diverse base categories (which may overlap with novel concepts) will ultimately improve generalization to novel classes.
Foundational FSOD. Rather than explicitly filtering target classes from pre-training [12, 50, 3], practitioners will likely employ foundational vision-language models (VLMs) pre-trained on (potentially private) web-scale data [33, 14, 20, 35] and fine-tune them for their task. As VLMs’ pre-training datasets contain diverse concepts [30, 20] [^1] The CLIP [30] pre-training dataset contains 500,000 concepts, spanning many categories encountered in the real world., it is challenging to prevent concept leakage. Therefore, one may be hesitant to exploit VLMs for FSOD. However, the performance of foundational models is undeniable; state-of-the-art VLMs like GroundingDINO already dominate all leading FSOD methods on COCO (48.3 vs. 33.1 AP) without fine-tuning (cf. Table 1).
Multi-Modal Concept Alignment. Does the strong zero-shot performance of VLMs imply that few-shot detection is no longer an interesting problem? No! We find that the target class name is often an insufficient description of the target concept. For example, trailer in nuImages [1] is defined differently than trailer in web-scale data. As shown in Fig. 2, even human annotators require few-shot instructions to identify subtle aspects of the target concept [2]. Interestingly, annotator instructions are naturally multi-modal, often including visual examples and textual descriptions. We advocate for a FSOD setup that uses similar visual and language cues for VLM concept alignment. We refer to our proposed setup as Foundational FSOD (cf. Fig. 1).
Federated Few-shot Learning. In order to effectively align VLM concepts with $K$-shot multi-modal “instructions”, we leverage a simple but evidently underappreciated observation: few-shot object detection datasets are actually federated datasets [11]. A federated dataset is a dataset comprised of smaller subsets, where each subset is exhaustively annotated for only a single category. For example, cars may or may not appear in the background of the $K$ images annotated with motorcycles (cf. Fig. 3). However, existing FSOD methods incorrectly assume that no cars are present in the background of non-car images. Inspired by prior work in learning with federated datasets [48] and weakly-supervised learning [15, 40, 39], we demonstrate that fine-tuning VLMs with federated losses consistently improves over zero-shot inference (cf. Tables 2, 4).
Contributions. We present three major contributions.
  1. 1.
    We modernize FSOD benchmarks by embracing vision-language foundation models that are pretrained on internet-scale data. We highlight the practical challenge of using multi-modal few-shot examples to define the target semantic concept (as shown in Fig. 2).
  2. 2.
    We point out that existing FSOD benchmarks are actually federated datasets, and present simple strategies for fine-tuning VLMs.
  3. 3.
    We conduct extensive experiments to ablate our design choices and demonstrate that our simple approach achieves state-of-the-art results on the LVIS and nuImages Foundational FSOD benchmarks.
2 Related Works
Few-Shot Object Detection aims to detect new categories with limited training data [16]. Recent work explores two primary approaches: meta-learning and transfer learning. Meta-learning-based methods focus on acquiring generalizable features from a set of base classes, which can then be applied to identify novel classes. For example, [13] proposes a technique that re-weights features from base classes to predict novel classes. [43] proposes a framework addressing both few-shot object detection and few-shot viewpoint estimation. [6] introduces a general FSOD network that learns a matching metric between image pairs, while [41] enhances object features using a universal prototype. More recently, [45] proposes a generative approach that is robust to noisy object proposals for novel classes. In contrast, transfer learning involves partially freezing network weights pretrained on a base dataset to improve a model’s ability to generalize to novel classes with limited data. Transfer learning approaches often follow a two-stage fine-tuning strategy: first train on base classes and then fine-tune the box classifier and regressor with K-shots from novel classes. This strategy has historically outperformed meta-learning approaches [38]. Recent work has primarily focused on improving classification performance. [36] utilizes a contrastive proposal encoding loss to encourage instance-level intra-class compactness and inter-class variance. Similarly, [19] applies a class margin loss to balance inter and intra-class margins. Our approach leverages transfer-learning by fine-tuning vision-language models (VLMs) pre-trained on large-scale datasets.
Vision Language Models are trained on a large-scale collection of weakly-supervised image-text pairs collected from the web. These models embed images and text into a shared space, enabling open-vocabulary detection. Early works adapt VLMs for object detection by either distilling the model’s predictions for specific image regions [10, 9] or directly incorporating detection components into frozen [17] or fine-tuned [27, 28, 4] encoders. In contrast, RegionCLIP [47] employs a multi-stage training approach, which involves generating pseudo-labels from captioning data, conducting region-text contrastive pre-training, and fine-tuning on detection data. GLIP [20] uses a single text query for the entire image and frames detection as a phrase grounding problem. Detic [49] addresses long-tail detection performance by leveraging image-level supervision. In the context of open-vocabulary detection, there may be some overlap between categories seen during training and testing. We use the term “zero-shot inference” to signify that a model has never been trained on the target dataset.
Federated Datasets are constructed by combining smaller datasets, each resembling a conventional object detection dataset for a single category [11]. Each of these smaller datasets ensures exhaustive annotations for a specific category. Images within each smaller dataset may overlap, resulting in some images with exhaustive annotations for multiple categories. Importantly, since exhaustive annotations for a particular category are only guaranteed within each small dataset, most images are sparsely annotated. As a result, naively training models with federated datasets leads to much sparser gradients [37]. To address this challenge, CenterNet2 [48] introduced FedLoss, a simple modification of cross-entropy loss which randomly samples a subset of negative categories for each image. We adopt FedLoss for FSOD, achieving consistent improvements over zero-shot inference.
Figure 3: Identifying Negative Pseudo-Labels. On the left, we visualize the standard $K$-shot detection setup, which we argue is actually a federated dataset [11] where one is given multiple mini-datasets of $K$ images. In this case, we visualize two $K=1$ datasets of bus and motorcycle. Importantly each mini-dataset does not provide information about the presence of other objects. Previous FSOD methods apparently ignore this fact, and instead assume the collective set of few-shot images are fully annotated across all object classes. This will likely produce many incorrect negative labels – e.g., all unlabeled cars in the background of the motorcycle mini-dataset will be incorrectly treated as negative cars. We use a ✓  to denote that a given image will be treated as a negative example of a given class by the learner and a ✗  to denote that a given image will be ignored when learning a given class. We color such negative labels as green when correct and red when incorrect. Naive FSOD approaches learn about all classes from all images, which results in many incorrect negative labels (shown in red on the left). Instead, we embrace the partially-labeled nature of the data and exploit tools from weakly-supervised learning, such as the use of psuedo-labels predicted by a teacher. We train (or rather, fine-tune) initial detectors on only the appropriate mini-dataset and use thresholded psuedo-detections (shown as cyan boxes on the right) to find images that can be confidently treated as (pseudo) negatives, which results in much fewer mistakes (shown in red on the right). This in turns produces improved performance. We also attempted to learn from psuedo positive labels, but found these to be less reliable. We include pseudo-code in the supplement.
Weakly Supervised Learning techniques are especially popular when learning in data-constrained settings, leveraging large-scale noisy annotations to improve model performance. Prior works [15] learn from noisy annotations using negative labels for classification. Further, [18, 44] leverage pseudo-labels for self-training. We modify our fine-tuning approach to leverage negatives derived from pseudo-labels to improve FSOD performance.
3 FSOD with Vision-Language Models
As shown in Fig 1, our proposed Foundational FSOD benchmark uses vision-language models (VLMs) pre-trained on diverse, large-scale datasets prior to fine-tuning on $K$ examples for each target class. We contrast our proposed setup with standard benchmarks, demonstrate that FSOD benchmarks are actually federated datasets, and present simple strategies for fine-tuning VLMs below.
3.1 Foundational FSOD Benchmark
Existing FSOD benchmarks repurpose well-established datasets like PASCAL VOC [5] and COCO [21] by partitioning them into base and novel classes for pre-training and fine-tuning, respectively. For COCO, the 60 categories disjoint with PASCAL VOC are used as base classes and the remaining 20 are used as novel classes [38]. However, this setup is artificial and does not reflect how FSOD is deployed in practice. First, it requires FSOD methods to detect common categories such as car and person by assuming they have only few examples. Importantly, VLMs like GroundingDINO [23] can already detect common categories with high accuracy on COCO without fine-tuning (cf. Table 1). Therefore, we focus on benchmarking Foundational FSOD on more realistic and challenging datasets like LVIS [11] and nuImages [1]. In addition, existing FSOD benchmarks require that datasets are partitioned into base and novel classes, which is not feasible for large-scale (often private) foundational datasets. For example, although CLIP’s [30] model weights are publicly available, its pre-training dataset is not. Instead, FSOD methods should only fine-tune VLMs on $K$-shot annotations for $C$ target classes, and also evaluate performance on these $C$ classes.
3.2 Few-Shot Multi-Modal Concept Alignment
Although VLMs achieve strong zero-shot performance on common classes, they struggle when the target class is different from vision-language concepts encountered in web-scale training (cf. Fig. 2). For example, nuImages [1] defines trailer as independent from the truck cab. However, Detic jointly detects the truck cab and its trailer together. This fine-grained distinction is provided to human annotators with visual examples and textual descriptions. Similarly, we provide few-shot multi-modal examples to align Detic to each target class.
We start by naively fine-tuning Detic [49] on $K$-shot vision-language examples. We ablate the impact of freezing different parts of Detic in Table 5 and find that freezing the backbone, RPN, and classifier head with CLIP embeddings yields the best performance (cf. Figure 4). However, due to the sparsity of annotations in the few-shot examples, we posit that Detic will receive sparser gradients [37]. Importantly, this degrades detection accuracy because all unannotated objects are treated as negatives. Therefore, we explore three strategies for handling negatives as described below.
3.3 FSOD Benchmarks are Federated Datasets
Prior works follow the $K$-shot dataset creation process established by [38]. To construct a $K$-shot dataset, we select a target class $C$ and an image at random. If the total annotations for class $C$ in the image are less than or equal to $K$, we add the image to our dataset. We repeat this process for all classes until we have exactly $K$ annotations per class. Importantly, each image in the dataset is exhaustively annotated for a subset of all classes. Recall, a federated dataset is also comprised of images that are exhaustively annotated for a specific category. This suggests that we can leverage insights about federated datasets [11, 48] to train better few-shot object detectors.
Fine-Tuning with FedLoss. We fine-tune Detic with Federated Loss (FedLoss) [48] using a subset $S$ of classes $C$ for each training image. Specifically, we use a binary cross-entropy loss on all classes in $S$ and ignore classes outside of $S$ during training. $S$ is comprised of the ground-truth annotation class along with randomly sampled negative classes for each image. We sample these negative classes in proportion to their square-root frequency in the training set. We find that probablistically sampling negatives rather than labeling all unannotated classes as negatives improves fine-tuning results, reliably beating zero-shot performance. Importantly, although FedLoss has been explored in the context of long-tailed detection, applying it to FSOD provides considerable performance improvements, reaffirming that FSOD benchmarks are actually federated datasets.
Fine-Tuning with Inverse FedLoss. FedLoss samples common classes like car more frequently as negatives, hurting detection accuracy for long-tailed datasets like LVIS and nuImages. Instead, we propose Inverse FedLoss (InvFedLoss), a minor modification of FedLoss that samples negative categories in proportion to the inverse of their square root frequency. This ensures that we sample rare categories as negatives more frequently to better match the true data distribution. Leveraging this insight improves over FedLoss and naive fine-tuning.
Fine-Tuning with Pseudo-Negatives. Despite the effectiveness of InvFedLoss, probablistically sampling negatives using dataset-wide statistics is sub-optimal because it does not consider the content of each image. We can improve the accuracy of sampled negatives with pseudo-labels to determine which classes are likely not in a particular image. If the maximal score for any class prediction is less than a threshold, we consider this class to be a negative. Using image predictions to identify pseudo-negatives yields better results than simply using dataset-wide statistics. We find that this strategy works the best. We present pseudo-code in the supplement.
Table 1: VLM Zero-Shot Inference Is a Strong FSOD Baseline. Zero-shot inference with VLMs like GroundingDINO resoundingly outperform state-of-the-art FSOD methods on the COCO FSOD benchmark, motivating the need to re-frame FSOD to embrace foundation models.
Approach 30-shots
AP bAP nAP
FRCN-ft-full [46] 18.6 20.6 12.5
FRCN-BCE [46] 30.2 36.8 10.3
TFA w/ fc [38] 29.3 34.5 13.5
TFA w/cos [38] 29.9 35.3 13.6
MPSR [42] 17.1 18.1 14.1
Meta-RCNN [46] 7.8 7.1 9.1
FsDetView [43] 10.0 9.3 12.0
Retentive R-CNN [7] 32.9 39.3 13.8
DiGeo [25] 33.1 39.4 14.2
GroundingDINO (Zero-Shot) [23] 48.3 46.3 54.3
4 Experiments
We conduct extensive experiments to validate that zero-shot inference from VLMs significantly improves over state-of-the-art FSOD approaches, suggesting that existing benchmarks should be re-framed to include foundation models. Moreover, we demonstrate that using federated losses consistently improve fine-tuning performance on LVIS and nuImages. Lastly, we compare our performance against an (oracle) upper bound and ablate the impact of freezing different detector components on fine-tuning performance.
Datasets and Metrics. We repurpose two established datasets for Foundational FSOD:
  • LVIS [11] re-annotates COCO images using 1,230 fine-grained classes, which are divided into frequent, common and rare based on the cardinality of each class. Frequent and common classes are combined to form LVIS-base and is used for pre-training. Rare classes are used for LVIS-novel. Following [38, 25], we benchmark with LVIS v0.5 on publicly released data splits and report performance averaged across 3 splits for frequent, common, and rare groups ($AP_{f},AP_{c},AP_{r}$) on the LVIS val-set.
  • nuImages [1] annotates 18 classes, which are divided into groups with many, medium, and few examples [29, 26] and report AP for each cohort. Although this dataset is not traditionally used for FSOD, nuImages’ open-world categories like debris and pushable-pullable make it particularly challenging (even for VLMs), and is a realistic benchmark for Foundational FSOD. Unlike LVIS, nuImages is fully annotated, so we construct our Foundational FSOD benchmark using the procedure described in subsection $3.$ Since the specific $K$ examples can have a significant impact on the overall performance, we run each experiment over 10 random data splits and report the average.
4.1 Zero-Shot Inference Is A Strong FSOD Baseline
Table 2: LVIS Foundational FSOD Performance. Detic pre-trained only on LVIS-base outperforms specialized methods like TFA and DiGeo by $\sim$6 AP, without fine-tuning on rare classes. Importantly these performance improvements can be attributed to Detic’s CLIP-based classifier and demonstrates that concept leakage through language can be advantageous in data constrained settings. By leveraging our insight that FSOD benchmarks are actually federated datasets, we show that fine-tuning with pseudo-negatives improves over standard fine-tuning by 1.6 AP (30.0 vs. 31.6, ResNet-50 backbone) and $4.3AP_{r}$ (15.5 vs. 19.8, ResNet-50 backbone). Lastly, we note that simply swapping the ResNet-50 backbone to a Swin transformer significantly improves performance (without modifying training data). Notably, we find that Detic’s rare class performance improves over naive fine-tuning by $5.9AP_{r}$ (from 26.7 AP to 32.6 AP). All the methods in this table are pre-trained on LVIS-base.
Approach 10-shots
APAP $AP_{f}$ $AP_{c}$ $AP_{r}$
ResNet-50 Backbone
TFA w/ fc [38] 24.1 27.9 23.9 14.9
TFA w/ cos [38] 24.4 27.7 24.3 16.9
DiGeo [25] 24.9 28.5 24.6 17.3
Detic (Base Only) [49] 30.0 34.4 30.8 16.3
   + Fine-Tuning (Base + Novel) 30.0 33.2 31.9 15.5
   w/ FedLoss 30.8 33.9 32.7 17.4
   w/ InvFedloss 31.1 34.3 32.5 18.7
   w/ Pseudo-Negatives 31.6 34.8 32.8 19.8
Swin Backbone
GroundingDINO (Zero-Shot) [23]  [^2] GroundingDINO is pre-trained on Objects365, GoldG and Cap4M 27.4 32.7 23.3 18.1
Detic (Base Only) [49] 35.2 38.7 36.8 21.4
   + Fine-Tuning (Base + Novel) 35.9 37.1 37.8 26.7
   w/ FedLoss 36.5 36.7 38.3 30.4
   w/ InvFedloss 37.1 37.8 38.5 31.1
   w/ Pseudo-Negatives 37.2 37.7 38.2 32.6
We compare state-of-the-art FSOD methods with zero-shot inference from GroundingDINO [23] on COCO in Table 1. Surprisingly, GroundingDINO outperforms DiGeo [25] by 16.2% AP averaged across both base and novel categories despite never being trained on COCO images. GroundingDINO’s impressive performance is due to its large-scale multi-modal pre-training on Objects365 [34], GoldG [14] and Cap4M [20]. It is worth noting that GroundingDINO achieves higher AP on novel classes than base, suggesting that novel classes in existing benchmarks (e.g. car and person) are actually not rare in the real world. Therefore, FSOD benchmarks should be re-framed to reflect real-world applications.
4.2 Foundational FSOD with LVIS
In this section, we evaluate GroundingDINO [23] and Detic’s  [49] performance on the LVIS Foundational FSOD benchmark. We train Detic from scratch on LVIS-base with a ResNet-50 backbone for fair comparison with prior work [38, 25].
As shown in Table 2, Detic outperforms all recent FSOD baselines including DiGeo [25] by about $\sim$6 points on $AP_{c}$ and $AP_{f}$ and achieves $16.3$ $AP_{r}$ without ever seeing any rare class data (e.g by prompting Detic (Base Only) with the rare class names). Importantly, these performance improvements can be attributed to Detic’s CLIP-based classifier, which has been trained with significantly more than $K$ examples of each class. This highlights the role of language in data-constrained settings.
Further, fine-tuning Detic with pseudo-negatives improves overall performance by 1.6% (30.0 vs 31.6) over naive fine-tuning. Finally, we note that simply replacing the ResNet-50 backbone with a Swin-B transformer yields a sizeable 12.8 AP improvement for rare classes (19.8 vs. 32.6).
Figure 4: Multi-Modal Fine-Tuning Improves Concept Alignment. While VLMs demonstrate strong zero-shot capabilities, they struggle when the target class is different from concepts encountered in web-scale training. Specifically, we find that GroundingDINO [23] (left) and Detic [49] (center) struggle to detect open-world categories like pushable-pullable. Fine-tuning Detic (right) with federated losses improves VLM concept alignment. Note that red boxes represent ground-truth annotations and green boxes are predictions by their respective models.
Table 3: nuImages Foundational FSOD Performance. We repurpose nuImages for FSOD following the dataset creation process established by [38]. We group categories by frequency into cohorts with many, medium and few examples per class (according to the fully-annotated dataset) [29, 26]. We fine-tune Detic (pre-trained on LVIS, COCO and ImageNet-21K) on $K$ examples for each of the 18 nuImages classes. We observe that existing methods like TFA perform poorly on nuImages, achieving <3<3AP. However, by simply pre-training TFA on more data and leveraging language cues via CLIP classifier, we dramatically improve performance (from 2.0 AP to 14.8 AP, 5-shots). Interestingly, accuracy across cardinalities decreases despite all classes being trained with $K$ examples. This suggests that despite pre-training on web-scale datasets, VLMs still struggle to detect rare categories like strollers, pushable-pullable, and debris, highlighting the challenge of working with nuImages.
Approach Average Precision (AP)
All Many Medium Few
GroundingDINO (Zero-Shot) [23] 11.44 17.42 16.13 3.38
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
5-shots
   TFA [38] w/ COCO-base 1.33 2.78 1.43 0.23
   TFA [38] w/ LVIS-base 2.02 1.69 4.08 0.58
   TFA [38] w/ LVIS,IN-21K, 14.77 25.16 18.65 3.63
   COCO + CLIP Classifier
   Ours 15.94 28.47 19.53 3.50
10-shots
   TFA [38] w/ COCO-base 1.21 2.55 1.19 0.31
   TFA [38] w/ LVIS-base 2.27 2.05 4.51 0.58
   TFA [38] w/ LVIS,IN-21K, 15.53 26.01 19.93 3.88
   COCO + CLIP Classifier
   Ours 16.67 29.15 20.66 3.90
30-shots
   TFA [38] w/ COCO-base 1.14 2.81 0.84 0.23
   TFA [38] w/ LVIS-base 2.23 1.48 4.98 0.45
   TFA [38] w/ LVIS,IN-21K, 16.83 27.90 21.59 4.45
   COCO + CLIP Classifier
   Ours 17.87 30.32 22.35 4.70
4.3 Foundational FSOD with nuImages
In the context of foundational models, we argue that partitioning datasets into base and novel classes no longer makes sense. Instead, FSOD methods should only train on $K$-shot annotations for $C$ target classes, and also evaluate performance on these $C$ classes. We fine-tune Detic (pre-trained on LVIS, COCO and ImageNet-21K) on $K$ examples per class and highlight model performance in Table 3. As expected, detection accuracy improves as we add more training examples. Despite large-scale pre-training, we see low accuracy for classes with few examples, highlighting the difficulty of the nuImages dataset. High intra-class variance for categories such as debris makes it difficult to generalize given few examples. According to nuImages’ annotation instructions, debris can include anything that is too big to be safely driven over, including fallen tree branches and trash bags. Similarly pushable-pullable includes trash cans, luggage, dollies, wheel barrows, and shopping carts.
To contextualize our results, we evaluate TFA [38] on the nuImages Foundational FSOD benchmark. We train two variants of TFA trained on COCO-base and LVIS-base and fine-tune both models on $K$ examples of the nuImages classes. Surprisingly, both variants of TFA achieve less than 33 AP (cf. Table 3). We posit that this is largely due to poor classification performance. Since both LVIS and COCO classes do not significantly overlap with nuImages classes, learning a classifier from few examples is extremely difficult. However, we find that simply re-training TFA with a frozen CLIP-based classifier (similar to Detic) dramatically increases performance, reiterating the utility of language and web-scale pre-training in data-constrained settings.
4.4 Oracle Performance Analysis
We compute the upper bound performance when given access to ground-truth negatives and exhaustive annotations for the few-shot data split. Recall, nuImages is exhaustively annotated, but is repurposed for Foundational FSOD.
To compute the set of ground-truth negatives for each image, we use exhaustive ground-truth annotations to determine which categories are not present. Note that this information doesn’t exist in LVIS because its ground-truth is sparsely annotated. Training with ground-truth negatives provides an upper bound on our pseudo-negatives experiment. Next, we train using exhaustive ground-truth annotations to provide an upper bound for the specific set of images used during training. In addition, this experiment highlights the performance gap between having exhaustive negatives and exhaustive annotations.
Table 4 shows that using pseudo-negatives nearly matches the true negative upper bound (16.67 AP vs 16.99 AP). This demonstrates that we are able to reliably estimate negatives in an image, alleviating the problem of learning with sparse annotations. Training with exhaustive annotations yields significantly better results for many and medium classes. This is unsurprising because the 10-shot FSOD benchmark includes 10 car annotations, while the exhaustively annotated set includes over 550 car annotations!
Despite strong performance on classes with many and medium, the upper bound for classes with few examples remains low (4.21 AP and 3.93 AP). We posit that it is very hard to capture the correct semantics of nuImages’ rare categories only using $K$-shots. We observe similar trends for the 5 and 30-shot cases and present further analysis in the supplement. Given the success of training with pseudo-negatives, a natural next-step is to train with pseudo-positives. Our preliminary results suggest that incorporating pseudo-positives does not provide significant improvement over simply training with pseudo-negatives. We posit that training with incorrect pseudo-positives may incur a higher penalty than training with incorrect pseudo-negatives. This is a promising direction for future work.
Table 4: Analysis of nuImages Upper Bound Performance. We compare the accuracy of our proposed approach against upper bounds computed for the FSOD task. Our pseudo-negatives strategy approaches the performance of using ground-truth negatives, demonstrating that pesudo-labels can provide a reliable signal about negatives, especially across classes with many and medium examples. The performance gap between our best method and exhaustive annotations can be attributed to the large number of additional annotations, particularly for classes with many and medium examples. Compared to the baseline (14.3 AP), our approach (16.7 AP) closes the gap to the (18.5 AP) upper-bound by over 50%.
Approach 10 Shots: Average Precision (AP)
All Many Medium Few
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
   + Fine-Tuning 15.53 26.01 19.93 3.88
   w/ FedLoss 15.57 27.20 20.09 2.89
   w/ Inverse FedLoss 15.89 27.56 20.19 3.42
   w/ Pseudo-Negatives 16.67 29.15 20.66 3.90
   w/ True Negatives (Oracle) 16.99 29.60 20.93 4.21
   w/ Exhaustive Annotations (Oracle) 18.51 33.51 22.35 3.93
4.5 Fine-Tuning Ablation
We explore different fine-tuning strategies for training Detic with few-shot annotations. We broadly divide Detic’s architecture into four components: Backbone, Region Proposal Network (RPN), Box Regressor, and Classifier. We ablate the impact of freezing different components and present results in Table 5.
Intuitively, as we have limited training data, we attempt to fine-tune a minimal number of parameters. Initializing and freezing the classifier head with CLIP embeddings corresponding to class names provides the most significant improvement. Prior works that fine-tune vision-only models have no notion of language embeddings and therefore must train classifiers from scratch. In contrast, Detic can represent class names using CLIP embeddings and can more easily adapt to novel categories with few examples. We find that freezing the backbone, RPN and classifier head with CLIP embeddings, and training the classifier projection layer and box regressor performs the best. Importantly, as Detic has been trained on large-scale datasets, its RPN can easily localize novel objects without fine-tuning.
[…]
Detic Components 10 Shots: Average Precision (AP)
Backbone RPN Box Regressor Classifier All Many Medium Few
- - - - 12.11 19.41 18.44 0.87
[…]
$faSnowflake \faSnowflake -$ 10.66 18.54 15.53 0.56
$faSnowflake \faSnowflake$ CLIP 15.31 26.83 19.58 2.89
4.6 Limitations and Future Work
Despite using VLMs pre-trained on large-scale datasets, we find that performance for rare categories (as defined by the cardinality of each class in the original dataset) is considerably lower than for common classes. We posit that VLMs are pre-trained with imbalanced data which includes many examples of common categories like truck but few examples of rare categories like stroller. Our work does not explicitly improve detection performance on the rare classes. Interestingly, since VLMs like Detic [49], GLIP [20], and GroundingDINO [23] are trained with different data sources, each model has dramatically different zero-shot performance on novel categories like stroller. Ensembling predictions from different VLMs may yield better detection accuracy for rare categories. In addition, although our work motivates the use of rich textual descriptions for multi-modal alignment, our approach only uses class names as text features. We hope future work can address the above limitations.
5 Conclusion
We revisit few-shot object detection (FSOD) with vision-language models (VLMs) and find that zero-shot inference from web-scale VLMs significantly outperforms leading FSOD methods. However, such foundational models do not fully address few shot recognition because of the concept alignment problem; particular concepts in target applications may be different than their use on web-scale datasets. Just as human annotators require concept alignment via multimodal text and visual examples, we argue that VLMs should be aligned with such few-shot data, formalizing the problem of Foundational FSOD. We also point out that existing FSOD benchmarks are actually federated datasets, and demonstrate that federated losses improve Foundational FSOD performance, approaching the oracle upper bound where few-shot images are fully annotated. Our analysis suggests that future few-shot (or “foundational concept alignment”) benchmarks may benefit from assuming images are fully annotated, since the cost of annotating a small set of $K$ images is manageable in practice.
Appendix A Implementation Details
Pseudo-Negative Federated Loss. First, we sample an image at random. We start by running a pre-trained detector on the image to generate predictions. We partition the predictions into two sets based on a confidence threshold; all predictions above the threshold are pseudo-positives (pseudo_pos) and all predictions below the threshold are ignored. Next, we compute negative classes (neg_classes) based on the pseudo-positives. Specifically, all classes not included in the psuedo-positive predictions are considered negative classes. Following the federated loss described in [48] (modified to use neg_classes instead of randomly sampled negative classes), we compute the binary cross entropy loss for the negative (neg_classes) and ground truth classes. See Algorithm 1 for pseudo-code.
# Inputs
# img: Randomly Sampled Image
# all_classes: All Classes in Dataset
# gt: Ground Truth Annotations for img
# gt_classes: List of Classes in gt
# Outputs
# loss: Psuedo-Negative Federated Loss
# Functions
# filter: Returns All Predictions w/
# Confidence > Threshold
# get_neg: Returns List of Classes Not
# In Pseudo-Positives
# or: Set Union Operation
# BCE: Binary Cross Entropy Loss
#Step 1: Compute Predictions
# and Filter by Confidence
pred = Detector(img) # predictions
pseudo_pos = filter(pred, thresh=0.2)
#Step 2: Get Pseudo-Negatives for Image
neg_classes = get_neg(pseudo_pos, all_classes)
select_classes = or(neg_classes, gt_classes)
#Step 3: Compute Deterministic Federated Loss
# w/ Pseudo-Negatives
loss = 0
for cls in select_classes:
pred_cls = pred[cls] #predictions for cls
gt_cls = gt[cls] #ground-truth for cls
loss += BCE(pred_cls, gt_cls)
return loss
Algorithm 1 Psuedo-Negative Federated Loss
LVIS v0.5 Experiment Details. We select Detic with a Resnet-50 backbone for fair comparison with prior work. We pre-train Detic on LVIS-base for $90k$ iterations with a batch size of $32$ using an AdamW optimizer and a learning rate of 2e32e-3. All images are resized to $640\times 640$ and we also enable Repeat Factor Sampling [11].
Following [38], we sample up to $10$ shots for each class in LVIS (since all classes may not have 10 examples). We use a batch size of $32$, learning rate of 2.5e52.5e-5 for $46k$ iterations. We do not use Repeat Factor Sampling for fine-tuning. We sample $50$ categories for each training image, i.e $|S|=50$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
nuImages Experiment Details. We select Detic with a Swin-T backbone pre-trained on LVIS+COCO and ImageNet-21k data. We use an image size of $1600\times 900$, batch size of 88 and an AdamW optimizer with learning rate of 3.75e63.75e-6. We fine-tune this model for $8000$ iterations on nuImages. We sample 66 categories for each training image, i.e $|S|=6$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
The original Detic implementation samples $|S|$ categories per batch, rather than per image. This works for LVIS, which has 1200 categories. However, nuImages only has $18$ classes, which required re-implementing the negative category sampling step for each image.
Appendix B Analysis of Iconic Few-Shot Images
The specific examples used during few-shot fine-tuning significantly impacts target class performance [38]. However, prior work constructs few-shot splits by randomly sampling $K$ examples per class. In contrast, when creating annotator instructions, selecting the right examples to “align” human annotators [2] to subtle aspects of the target concept is carefully considered. To more closely match VLM concept alignment with human annotator alignment, we design a simple algorithm to construct the best $K$-shot split for fine-tuning. This allows us to understand which examples are most informative and measure an upper bound in performance.
We construct our best split by picking examples corresponding to the best class-wise performance, based on the evaluation of each split on a held-out validation set. For instance, out of $10$ random splits for the 55-shot task, one may pick car examples from split 11, bicycle from split 44 and debris from split 88. In Table 6, we observe that the best-split performance is always better than its random counterpart. As expected, the choice of examples in 55-shot case is more important than the $30$-shot case ($1.63$ AP difference for 55-shot vs $0.08$ AP for $30$-shots). We visualize the difference in the splits for the stroller class in nuImages (cf. Figure 5). Unsurprisingly, iconic examples are large and unoccluded.
Figure 5: Visualizing Random and Best Split. In the top row, we visualize the 5-shot training examples of strollers from a random split. Similarly, we visualize the 5-shot training examples from the best split in the bottom row. We observe that strollers in the random split are often occluded, small in size and are blurry, making few-shot learning harder. On the other hand, the best split examples are larger, have better visual quality and are relatively un-occluded. This visual difference directly translates into better few-shot performance. We achieve $\mathbf{13.09}$ Stroller AP for the random split and $\mathbf{18.54}$ Stroller AP for the best split. We perform a more comprehensive evaluation in Table 6.
Table 6: Random Split vs ‘Best” Split. We construct the “best” split by selecting per-class few-shot examples that lead to the highest performance on a held out set. Unsurprisingly, the best split performs better than any random split, especially for very limited data settings (e.g. 55-shot detection). This evaluation setting closely mimics how human annotators are “aligned” to target concepts, since annotator guides are constructed using hand-picked iconic visual examples.
Approach Average Precision (AP)
All Many Medium Few
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
5-shots
   Ours (Random Split) 15.94 28.47 19.53 3.50
   Ours (Best Split) 17.57 29.99 20.70 4.92
10-shots
   Ours (Random Split) 16.67 29.15 20.66 3.90
   Ours (Best Split) 17.51 29.98 21.60 4.58
30-shots
   Ours (Random Split) 17.87 30.32 22.35 4.70
   Ours (Best Split) 17.95 29.50 22.69 4.85
Appendix C nuImages Foundational FSOD Performance for K={5, 30} Shots
We observe similar trends for $K=\{5,30\}$ to the 10-shot performance reported in the main paper. Note that the reported results are averaged over $10$ seeds to reduce variance. For $K=5$ shots, we note that naive fine-tuning barely improves the overall AP (cf. Table  7). Incorporating pseudo-negatives in this data-scarce setting closes the gap to the true negatives upper bound ($15.94$ vs $16.15$ AP). Even with exhaustive annotations, we only improve by $\sim 3.3$ AP ($14.26$ vs $17.59$ AP) points over the zero-shot results.
Naive fine-tuning improves over zero shot results by $\sim 2.5$ AP ($16.83$ vs $14.26$ AP) in the $K=30$ shot setting. In addition, we see diminishing benefits for using FedLoss variants (cf. Table 8). Nevertheless, pseudo-negatives improve over naive fine-tuning by 11 AP ($16.83$ vs $17.87$ AP). Unsurprisingly, performance for classes with many examples increases the most ($19.40$ AP) when training with exhaustive annotations. Importantly, training with exhaustive annotations uses orders of magnitude more examples (1300 examples of car) than with few-shot annotations (30 examples of car).
Table 7: Analysis of 55-shot Performance on nuImages. Fine-tuning with pseudo-negatives improves over naive fine-tuning. Importantly, our pseudo-negative approach nearly matches the upper bound with true negatives ($15.94$ AP vs $16.15$ AP). This suggests that our proposed approach is particularly effective in data-constrained settings.
Approach 5 shots: Average Precision (AP)
All Many Medium Few
Detic [49] Zero-Shot 14.26 27.28 16.88 2.36
   + Fine-tuning             14.77 25.16 18.65 3.63
   w/ FedLoss 15.03 26.41 19.07 3.03
   w/ Inverse FedLoss 15.21 26.77 19.01 3.29
   w/ Pseudo-Negatives 15.94 28.47 19.53 3.50
   w/ True Negatives (Oracle) 16.15 29.00 19.52 3.70
   w/ Exhaustive Annotations (Oracle) 17.59 32.66 20.93 3.43
Table 8: Analysis of $30$-shot performance on nuImages Naive fine-tuning provides a considerable improvement over zero-shot inference due to a greater number of examples per-class. In constrast to 5-shot results, FedLoss variants provide limited improvement over naive fine-tuning. We find that fine-tuning with pseudo-negatives provides a 1% improvement overall.
Approach 30 shots: Average Precision (AP)
All Many Medium Few
Detic [49] Zero-Shot 14.26 27.28 16.88 2.36
   + Fine-tuning             16.83 27.90 21.59 4.45
   w/ FedLoss 16.45 28.88 21.14 3.02
   w/ Inverse FedLoss 16.88 29.19 21.41 3.77
   w/ Pseudo-Negatives 17.87 30.32 22.35 4.70
   w/ True Negatives (Oracle) 18.19 30.61 22.66 5.11
   w/ Exhaustive Annotations (Oracle) 19.40 34.24 23.75 4.48
Table 9: RegionCLIP Experiments. RegionCLIP zero-shot inference performs much worse than Detic. While fine-tuning improves RegionCLIP’s performance, it still lags far behind Detic. We posit that this performance difference can be attributed to Detic’s box-supervised pre-training and use of language cues from CLIP embeddings.
Approach Average Precision (AP)
All Many Medium Few
RegionCLIP (Zero-Shot) [47] 2.34 3.33 3.87 0.22
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
5-shots
   RegionCLIP (Fine-Tuning) [47] 3.61 6.20 5.14 0.26
   Detic (Fine-Tuning) [49] 14.50 24.09 18.53 3.70
10-shots
   RegionCLIP (Fine-Tuning) [47] 3.58 6.10 5.16 0.24
   Detic (Fine-Tuning) [49] 15.28 26.93 19.89 3.27
30-shots
   RegionCLIP (Fine-Tuning) [47] 3.57 6.13 5.10 0.22
   Detic (Fine-Tuning) [49] 16.65 27.45 21.51 4.02
Appendix D RegionCLIP Experiments
In this section, we evaluate the importance of using box supervised data in pre-training. Unlike Detic, which trains on box-supervised data from LVIS, COCO and image-text data from ImageNet21k, RegionCLIP[47] only pre-trains on image-text pairs from the Conceptual Captions (CC3M) dataset [35].
We report RegionCLIP’s zero-shot and fine-tuning performance on nuImages averaged over 33 random splits in Table 9. Detic zero-shot outperforms RegionCLIP zero-shot by $\sim 12$ AP ($14.26$ vs $2.34$). While fine-tuning RegionCLIP improves overall performance, Detic achieves higher accuracy for $K=\{5,10,30\}$ shots. This highlights the importance of supervision type (e.g. box-supervised data) and data scale used for pre-training.
Next, we conduct further analysis to diagnose why RegionCLIP zero-shot inference performs so poorly on nuImages (Table 10). RegionCLIP relies on an RPN pre-trained on box-supervised data like LVIS-base to extract regions for pre-training. Notably, RegionCLIP (w/ LVIS-RPN: $2.34$ AP) suffers from poor foreground-vs-background classification compared to Detic. We validate this hypothesis by evaluating RegionCLIP (w/ GT-RPN) to measure classification performance. Surprisingly, RegionCLIP achieves significantly higher accuracy ($26.44$ AP), confirming that RegionCLIP struggles to distinguish between foreground and background in nuImages. This observation highlights the challenge of working with nuImages categories, further motivating our Foundational FSOD benchmark.
Lastly, we evaluate RegionCLIP’s performance with Detic-RPN. Notably, we observe that the performance improves over RegionCLIP w/ LVIS-RPN demonstrating that reducing the number of false positive proposals yields better performance. Furthermore, we filter out low confidence Detic proposals , i.e <0.5<0.5 objectness score (w/ Detic-RPN, 0.5) and find that this doubles RegionCLIP’s zero-shot performance to $7.64$ AP.
Table 10: Diagnosing RegionCLIP’s Poor Zero-Shot Performance. RegionCLIP’s zero-shot performance lags far behind Detic. Using RegionCLIP’s classifier on ground-truth region proposals yields high performance, suggesting that RegionCLIP struggles to accuratly distinguish between foreground-vs-background.
Approach Average Precision (AP)
All Many Medium Few
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
GroundingDINO (Zero-Shot) [23] 11.44 17.42 16.13 3.38
RegionCLIP (Zero-Shot) w/ LVIS-RPN [47] 2.34 3.33 3.87 0.22
RegionCLIP (Zero-Shot) w/ Detic-RPN [47] 3.79 6.68 3.91 1.12
RegionCLIP (Zero-Shot) w/ Detic-RPN, 0.5 [47] 7.64 12.81 9.57 1.88
RegionCLIP (Zero-Shot) w/ GT-RPN [47] 26.44 45.33 31.83 3.92
Appendix E NuImages Per-Split Breakdown
We provide per-split breakdowns for the nuImage experiments in Tables 11, 12, and 13.
Table 11: 5-shot nuImages Per-Split Performance
Approach Split 1 Split 2 Split 3 Split 4 Split 5
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-tuning 15.07 24.89 18.82 4.58 13.27 20.83 17.54 3.58 15.15 26.55 19.23 2.93 14.82 25.47 19.10 3.20 14.76 23.96 19.45 3.69
FedLoss 15.27 26.59 19.25 3.41 15.08 27.23 18.59 2.80 15.22 26.69 19.34 2.97 14.72 25.42 19.44 2.65 15.03 25.87 19.39 3.01
InvFedLoss 15.36 28.54 18.59 2.77 15.07 28.17 18.04 2.71 15.40 27.91 18.96 2.94 15.04 26.97 18.95 2.80 15.55 28.28 18.93 3.12
Pseudo-Negatives 15.87 28.59 19.06 3.67 15.88 28.70 19.28 3.30 16.12 27.94 20.01 3.84 15.76 28.00 19.65 3.37 16.21 28.43 19.90 3.80
True Negatives (Oracle) 16.27 28.55 19.35 4.56 15.92 29.03 19.24 3.22 15.97 28.81 19.42 3.29 15.92 28.61 19.57 3.41 16.44 28.91 20.1 3.92
Exh. Annotations (Oracle) 17.59 32.63 20.3 3.97 17.71 32.91 21.23 3.31 17.13 32.03 20.42 3.02 17.81 32.86 21.51 3.41 17.72 32.88 20.99 3.48
Approach Split 6 Split 7 Split 8 Split 9 Split 10
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 14.94 24.77 19.37 3.84 14.58 25.02 18.26 3.54 15.15 26.56 19.14 3.38 15.39 26.42 18.90 4.20 14.58 27.13 16.73 3.38
FedLoss 15.27 26.30 19.82 3.06 14.88 27.23 18.41 2.59 14.85 25.27 19.68 3.04 15.25 26.40 19.15 3.58 14.79 27.09 17.61 3.16
InvFedLoss 14.65 23.63 19.56 3.67 14.21 23.34 18.73 3.55 14.86 24.71 19.82 3.44 15.74 27.73 19.13 3.97 15.29 28.43 18.19 3.02
Pseudo-Negatives 16.30 27.66 20.89 3.82 15.82 28.69 19.00 3.47 15.82 27.93 20.11 3.28 15.73 29.08 18.95 3.04 15.87 29.67 18.46 3.37
True Negatives (Oracle) 16.46 29.5 20.38 3.43 15.95 29.54 18.94 3.26 16.23 28.4 20.14 4.18 16.52 29.4 19.77 4.12 15.8 29.3 18.31 3.6
Exh. Annotations (Oracle) 18.07 32.93 21.84 3.7 17.41 32.67 20.71 3.17 17.54 32.81 20.77 3.39 17.4 32.45 20.26 3.69 17.54 32.46 21.27 3.14
Table 12: 10-shot nuImages Per-Split Performance
Approach Split 1 Split 2 Split 3 Split 4 Split 5
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 15.77 26.63 20.13 3.87 15.31 26.85 19.46 3.03 15.64 27.32 20.08 2.91 15.78 25.16 20.79 4.44 15.93 27.34 20.28 3.60
FedLoss 15.66 27.74 20.05 2.83 15.53 25.71 20.84 3.07 15.62 28.49 19.69 2.43 15.63 27.14 20.31 2.91 15.72 25.22 21.37 3.54
InvFedLoss 15.96 27.22 20.51 3.66 15.80 26.05 20.88 3.54 15.85 28.71 19.97 2.62 16.04 27.43 20.49 3.68 16.26 27.05 21.22 3.80
Pseudo-Negatives 16.78 29.33 20.64 4.14 16.58 29.73 20.53 3.36 16.51 28.95 20.72 3.44 16.95 28.73 21.49 4.12 17.08 29.84 21.22 3.90
True Negatives (Oracle) 16.91 29.67 20.67 4.19 16.92 29.75 21.08 3.73 16.57 29.69 20.38 3.4 17.36 29.96 21.24 4.67 17.52 29.46 22.13 4.61
Exh. Annotations (Oracle) 18.03 33.17 21.33 3.77 18.68 33.98 22.52 3.88 18.3 33.2 22.37 3.45 18.44 33.47 22 4 19.08 33.71 23.36 4.42
Approach Split 6 Split 7 Split 8 Split 9 Split 10
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 15.22 23.86 21.10 3.65 15.00 22.86 19.67 5.07 15.67 27.42 19.67 3.55 15.25 25.67 19.61 3.60 15.73 27.02 18.52 5.10
FedLoss 15.79 26.93 21.08 2.82 15.57 27.22 19.76 3.17 15.47 27.98 19.46 2.73 15.32 27.02 19.92 2.44 15.40 28.60 18.45 2.92
InvFedLoss 16.07 27.49 21.09 3.15 15.93 28.80 19.43 3.31 15.69 28.38 19.45 3.04 15.29 25.94 20.12 2.94 15.99 28.50 18.72 4.43
Pseudo-Negatives 16.60 29.51 20.83 3.33 16.50 27.91 20.24 4.82 16.51 29.26 20.61 3.50 16.37 28.86 20.75 3.09 16.82 29.37 19.56 5.27
True Negatives (Oracle) 17.06 29.29 22.12 3.59 16.43 29.9 19.55 3.76 17.04 29.85 20.77 4.46 17.11 29.38 21.48 4.02 16.96 29.04 19.86 5.63
Exh. Annotations (Oracle) 18.62 33.65 23.09 3.45 18.75 33.83 22.2 4.51 17.99 33.11 21.45 3.78 18.33 33.18 22.47 3.51 18.94 33.77 22.67 4.55
Table 13: 30-shot nuImages Per-Split Performance
Approach Split 1 Split 2 Split 3 Split 4 Split 5
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 16.71 28.14 21.37 4.14 15.94 26.49 21.04 3.67 16.95 27.73 22.11 4.24 16.33 28.71 20.41 3.52 17.46 28.50 22.64 4.68
FedLoss 16.20 28.86 20.76 2.75 15.48 27.30 20.41 2.31 16.59 28.80 21.71 2.82 16.62 28.85 21.45 3.09 17.15 29.59 21.94 3.53
InvFedLoss 16.47 29.26 20.70 3.28 16.09 27.10 21.52 3.09 16.78 29.52 21.18 3.40 16.97 29.25 21.39 3.90 17.57 29.84 22.14 4.35
Pseudo-Negatives 18.14 30.45 22.33 5.38 17.34 29.24 22.53 3.91 17.90 31.09 22.03 4.42 17.83 30.44 22.34 4.35 18.33 30.78 23.04 4.87
True Negatives (Oracle) 18.34 30.55 22.52 5.68 17.82 29.46 22.89 4.83 17.76 31.09 21.93 4.15 18.36 30.99 22.61 5.18 18.67 31.03 23.59 5.18
Exh. Annotations (Oracle) 19.48 34.22 23.24 5.12 19.43 34.39 23.81 4.47 19.33 34.3 23.97 3.98 19.21 34.18 23.44 4.15 19.6 34.2 24.22 4.62
Approach Split 6 Split 7 Split 8 Split 9 Split 10
All Many Med Few All Many Med Few All Many Med Few All Many Med Few All Many Med Few
Zero-Shot 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36 14.26 27.28 16.88 2.36
Fine-Tuning 16.66 28.51 21.51 3.59 17.30 28.02 21.89 5.53 17.19 27.68 22.42 5.01 17.05 28.11 21.84 4.56 16.68 27.12 20.67 5.57
FedLoss 17.01 28.95 22.57 3.08 16.22 29.48 19.98 3.14 16.46 29.15 21.14 3.03 16.47 29.46 20.89 2.84 16.30 28.33 20.56 3.59
InvFedLoss 17.03 29.64 22.03 3.16 17.36 29.69 21.66 4.60 17.05 29.18 21.90 4.02 16.65 29.73 20.88 3.18 16.82 28.65 20.66 4.77
Pseudo-Negatives 18.08 30.58 23.00 4.39 18.05 30.20 22.37 5.43 17.61 30.47 22.01 4.33 17.87 30.13 22.50 4.64 17.56 29.77 21.34 5.29
True Negatives (Oracle) 17.79 30.57 22.92 3.71 18.44 30.47 22.66 6.07 18.3 30.83 22.72 5.34 18.39 30.74 23.02 5.08 18.07 30.35 21.8 5.88
Exh. Annotations (Oracle) 19.32 34.21 24.33 3.71 19.47 34.47 23.38 4.91 19.4 34.17 23.7 4.66 19.47 34.12 23.94 4.57 19.34 34.17 23.4 4.65